OpenStudy (anonymous):

Martin earned the following scores on his last five tests. 98, 78, 84, 75, 91 What is the interquartile range of his scores? 23 13 16 18 Can someone tell me how to do this? Thnx!

OpenStudy (anonymous):

You know how to do ranges, yes?

OpenStudy (notamathgenius):

98-75=x

OpenStudy (anonymous):

yup, i can do ranges,

OpenStudy (anonymous):

Then you can solve this question, no?

OpenStudy (anonymous):

no, what is an interquartile?

OpenStudy (notamathgenius):

It just means the range the answer is 23

OpenStudy (anonymous):

The answer is not 23.

OpenStudy (notamathgenius):

really hmmm wow my math is getting hiddeous

OpenStudy (anonymous):

:(

OpenStudy (kropot72):

The interquartile range is the difference between the upper and lower quartiles of a set of numbers. As an example the numbers 3 4 4 4 6 6 6 6 7 7 8 9 have upper quartile 7, lower quartile 4; so the interquartile range is 7 - 4 = 3

OpenStudy (anonymous):

I thought that was just range?

OpenStudy (anonymous):

when i do it that way, i get 23, and 23 is wrong.

OpenStudy (kropot72):

A quartile is the value of a variable below which one quarter (lower quartile) or three quarters (upper quartile) of a distribution lie. the median is the second quartile. Ordering the data in your question we get 75 78 84 91 98 Do you know which value is the median?

OpenStudy (kropot72):

@lillybeth123 Do you know how to find the median?

OpenStudy (anonymous):

yup i think it is 84.

OpenStudy (kropot72):

Good! You are correct. So if the median is not included we get two subsets {75, 78} and {91, 98} The quartiles are the medians of these two subsets which in this case need to be the arithmetic means of the two numbers. lower quartile = (75 + 78)/2 = ? upper quartile = (91 + 98)/2 = ? Can you calculate these two values?

OpenStudy (anonymous):

yup, 76.5 and 94.4

OpenStudy (kropot72):

Nearly right! The upper quartile is 94.5. Now to find the interquartile range of the scores just subtract the lower quartile from the upper quartile; 94.5 - 76.5 = ?

OpenStudy (anonymous):

oh, ok, 94.5, one minute for ur other problem.

OpenStudy (anonymous):

18.

OpenStudy (kropot72):

Good work! The correct answer choice is 18.

OpenStudy (anonymous):

ok, so i just do the same thing for my problem?

OpenStudy (kropot72):

Yes, 18 is the correct answer for the problem that you posted.

OpenStudy (anonymous):

Thanks!!!!! :))))

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Bye! :)